A coin is tossed three times. If Z is a random variable giving the number of heads obtained, (i) find the probability distribution of Z. (ii) what is the expected value of Z?


1
Expert's answer
2021-07-19T05:45:16-0400

(i)


"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 z &Outcomes & p(z) \\ \hline
 0 & TTT & 1/8 \\
 \hdashline
 1 & HTT, THT, TTH & 3/8 \\
 \hdashline
 2 & HHT, HTH, THH & 3/8 \\
 \hdashline
 3 & HHH & 1/8 \\
\end{array}"

"\begin{matrix}
 z & 0 & 1 & 2 & 3 \\
 p(z) & 1/8 & 3/8 & 3/8 & 1/8
\end{matrix}"

(ii)


"E[Z]=0(\dfrac{1}{8})+1(\dfrac{3}{8})+2(\dfrac{3}{8})+3(\dfrac{1}{8})=\dfrac{3}{2}=1.5"